# Gravitational waves in a minimal gravitational SME

### v1

@@ I Introduction
When gravity is treated as a perturbative field around flat spacetime, Lorentz–violating corrections appear in the quadratic sector of the action through operators of different mass dimensions. These contributions can be organized according to the operator dimension $d$ and separated based on their behavior under linearized coordinate transformations. In particular, invariance under the gauge transformation $h_{\mu\nu}\rightarrow h_{\mu\nu}+\partial_{\mu}\xi_{\nu}+\partial_{\nu}\xi_{\mu}$, with $h_{\mu\nu}$ denoting metric fluctuations about Minkowski spacetime and $\xi_{\mu}$ an infinitesimal vector field, serves as the criterion distinguishing diffeomorphism–preserving operators from those that explicitly break this symmetry. Within this classification, operators that respect diffeomorphism invariance arise only at mass dimension $d\geq 4$ Kostelecky:2017zob ; Kostelecky:2016kfm . The impact of such terms on gravitational–wave propagation and the corresponding observational bounds has been examined in detail in numerous studies, including Refs. Kostelecky:2016kfm ; Hou:2024xbv ; Dong:2023nau ; Zhu:2023rrx ; Mewes:2019dhj ; Kostelecky:2017zob . By contrast, if diffeomorphism symmetry is relaxed at the linearized level, Lorentz–violating operators of lower dimensionality become admissible, allowing contributions with $d=2$ and $d=3$ Kostelecky:2017zob .
@@ III Polarization properties of gravitational waves
On the other hand, the polarization content of gravitational waves encodes essential information about both their astrophysical origin and the fundamental symmetries governing gravity. In particular, departures from local Lorentz invariance can manifest directly in the polarization structure, making polarization observables a powerful probe of gravitational dynamics and wave propagation le2017theory ; liang2017polarizations ; zhang2018velocity ; hou2018polarizations ; mewes2019signals ; liang2022polarizations .
@@ V Phenomenological bounds
In the minimal gravity sector with operator dimension $d=4$, Lorentz-violating effects enter gravitational-wave propagation through a modified dispersion law. Focusing exclusively on the momentum-independent and nonbirefringent contribution, and keeping the speed of light explicit, the dispersion relation can be expressed as
$E\equiv p^{0}c=v\,|\vec{p}|\equiv v\,p,\qquad p\equiv|\vec{p}|,$ (49)
so that the propagation speed is
$v\equiv c\left(1-\mathring{k}^{(4)}_{(I)}\right).$ (50)
Because the coefficient $\mathring{k}^{(4)}_{(I)}$ does not depend on the wave momentum, the dispersion relation retains its linear form, implying that the phase and group velocities are identical
$v_{\mathrm{ph}}=\frac{E}{p}=v,\qquad v_{g}=\frac{\mathrm{d}E}{\mathrm{d}p}=v.$ (51)
In the minimal case with $d=4$, the theory predicts no birefringent behavior, implying that all polarization modes of gravitational waves travel at an identical speed. The resulting phenomenology is therefore limited to a uniform deviation of the propagation velocity from the speed of light, independent of polarization. Bounds on the coefficient $\mathring{k}^{(4)}_{(I)}$ are inferred using a phenomenological framework originally developed by Will Will:1997bb and later implemented in gravitational-wave analyses by the LIGO and Virgo
Collaborations abbott2016observation . Within this scheme, departures from luminal propagation are quantified by comparing the modified group velocity to that of an effective massive mode
$v_{g}^{(m)}\simeq c\left[1-\frac{1}{2}\left(\frac{m_{g}c^{2}}{E_{g}}\right)^{2}\right],$ (52)
with $m_{g}$ denotes the upper limit on the graviton mass, while $E_{g}$ represents the corresponding graviton energy. For the GW150914 signal, the characteristic peak frequency is measured to be $\nu\simeq 150,\mathrm{Hz}$ abbott2016observation . Identifying the graviton energy through the relation $E_{g}=h\nu$, with Planck’s constant given by $h=4.136\times 10^{-15},\mathrm{eV,s}$, one obtains
$E_{g}=h\nu\simeq 6.0\times 10^{-13}\penalty 10000\ \mathrm{eV}.$ (53)
Moreover, combining this estimate with the constraint $m_{g}<1.2\times 10^{-22},\mathrm{eV}/c^{2}$ leads to
$\Delta v_{g}\equiv c-v_{g}^{(m)}\;\lesssim\;\frac{c}{2}\left(\frac{m_{g}c^{2}}{E_{g}}\right)^{2}\;\lesssim\;2.0\times 10^{-20}\,c.$ (54)
Within this framework, departures from light-speed propagation are described by
$\Delta v\equiv|c-v|=c\left|\mathring{k}^{(4)}_{(I)}\right|.$ (55)
Imposing the condition $\Delta v\leq\Delta v_{g}$ then implies
$\left|\mathring{k}^{(4)}_{(I)}\right|\;\lesssim\;2.0\times 10^{-20}.$ (56)
An independent and significantly tighter bound is provided by multimessenger measurements. The almost simultaneous detection of the gravitational–wave event GW170817 and its electromagnetic counterpart, originating from a source at a distance of approximately $40,\mathrm{Mpc}$, constrains any deviation of the gravitational–wave speed from that of light to satisfy $|v-c|/c\lesssim 10^{-15}$ abbott2017gw170817 . Within the context of the present model, this observational limit implies
$\left|\mathring{k}^{(4)}_{(I)}\right|\;\lesssim\;10^{-15},$ (57)
resulting in a constraint that is more stringent than the one derived from the massive graviton benchmark.
@@ VII Acknowledgments
A. A. Araújo Filho is supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) and Fundação de Apoio à Pesquisa do Estado da Paraíba (FAPESQ), project numbers 150223/2025-0 and 1951/2025.

### v2

@@ I Introduction
When gravity is treated as a perturbative field around flat spacetime, Lorentz–violating corrections appear in the quadratic sector of the action through operators of different mass dimensions. These contributions can be organized according to the operator dimension $d$ and separated based on their behavior under linearized coordinate transformations. In particular, invariance under the gauge transformation $h_{\mu\nu}\rightarrow h_{\mu\nu}+\partial_{\mu}\xi_{\nu}+\partial_{\nu}\xi_{\mu}$, with $h_{\mu\nu}$ denoting metric fluctuations about Minkowski spacetime and $\xi_{\mu}$ an infinitesimal vector field, serves as the criterion distinguishing diffeomorphism–preserving operators from those that explicitly break this symmetry. Within this classification, operators that respect diffeomorphism invariance arise only at mass dimension  [^1] Here, it is worth mentioning that, in the context of the SME, the designation minimal refers to the restriction to power–counting renormalizable operators, namely those with mass dimension $d\leq 4$. In the gravitational sector, and in particular within the linearized framework adopted in this work, this restriction amounts to retaining only the $d=4$ Lorentz–violating operators in the quadratic action for metric perturbations, while excluding higher–dimension ($d>4$) nonminimal contributions. $d\geq 4$ Kostelecky:2017zob ; Kostelecky:2016kfm . The impact of such terms on gravitational–wave propagation and the corresponding observational bounds has been examined in detail in numerous studies, including Refs. Kostelecky:2016kfm ; Hou:2024xbv ; Dong:2023nau ; Zhu:2023rrx ; Mewes:2019dhj ; Kostelecky:2017zob ; Arzano:2016twc ; Kostelecky:2008ts ; Haegel:2022ymk ; ONeal-Ault:2021uwu . By contrast, if diffeomorphism symmetry is relaxed at the linearized level, Lorentz–violating operators of lower dimensionality become admissible, allowing contributions with $d=2$ and $d=3$ Kostelecky:2017zob .
@@ III Polarization properties of gravitational waves
On the other hand, the polarization content of gravitational waves encodes essential information about both their astrophysical origin and the fundamental symmetries governing gravity. In particular, departures from local Lorentz invariance can manifest directly in the polarization structure, making polarization observables a powerful probe of gravitational dynamics and wave propagation le2017theory ; liang2017polarizations ; zhang2018velocity ; hou2018polarizations ; mewes2019signals ; liang2022polarizations ; Nilsson:2023szw ; Bailey:2023lzy .
@@ V Phenomenological bounds
In the isotropic minimal gravitational SME restricted to the CPT–even $d=4$ sector, the tensor modes obey the strictly nondispersive dispersion relation
$\omega=\left(1-\mathring{k}^{(4)}_{(I)}\right)|\mathbf{p}|,$ (49)
so that the phase and group velocities coincide and are independent of frequency,
$v_{g}=1-\mathring{k}^{(4)}_{(I)}.$ (50)
The most stringent direct constraint on a frequency–independent deviation of the gravitational wave speed arises from the multimessenger observation of GW170817 and its electromagnetic counterpart GRB 170817A LIGOScientific:2017zic ; Abbott:2017oio ; Abbott:2017vwq . The near–coincident arrival of the gravitational and gamma–ray signals implies
$3\times 10^{-15}\;\lesssim\;\frac{v_{g}-1}{1}\;\lesssim\;7\times 10^{-16}.$ (51)
Using Eq. (50),
$\frac{v_{g}-1}{1}=-\mathring{k}^{(4)}_{(I)},$ (52)
which yields the conservative bound
$\left|\mathring{k}^{(4)}_{(I)}\right|\lesssim 3\times 10^{-15}.$ (53)
@@ VII Acknowledgments
A. A. Araújo Filho is supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) and Fundação de Apoio à Pesquisa do Estado da Paraíba (FAPESQ), project numbers 150223/2025-0 and 1951/2025. I. P. L. was partially supported by the National Council for Scientific and Technological Development - CNPq, grant 312547/2023-4. I. P. L. would like to acknowledge the networking support by the COST Action RQI (CA23115) and the COST Action FuSe (CA24101) supported by COST (European Cooperation in Science and Technology). N. H. would like to acknowledge networking support of the COST Action CA 22113 - Fundamental challenges in theoretical physics (Theory and Challenges), CA21106 - COSMIC WISPers in the Dark Universe: Theory, astrophysics and experiments (CosmicWISPers), CA21136 - Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse). The authors would like to acknowledge networking support by the COST Action BridgeQG (CA23130) supported by COST (European Cooperation in Science and Technology).